# Slope Consistency of Quasi-Maximum Likelihood

### v2

@@ Abstract
Although QMLE is generally inconsistent, logistic regression relying on the binary choice model (BCM) with logistic errors is widely used, especially in machine learning contexts with many covariates and high-dimensional slope coefficients. This paper revisits the slope consistency of QMLE for BCMs. Ruud (1983) introduced a set of conditions under which QMLE may yield a constant multiple of the slope coefficient of BCMs asymptotically. However, he did not fully establish slope consistency of QMLE, which requires the existence of a positive multiple of slope coefficient identified as an interior maximizer of the population QMLE likelihood function over an appropriately restricted parameter space. We fill this gap by providing a formal proof of slope consistency under the same set of conditions for any binary choice model identified as in Manski (1975, 1985). Our result implies that logistic regression yields a consistent estimate for the slope coefficient of BCMs under suitable conditions.
@@ 1 Introduction
Logistic regression is used extensively in machine learning to analyze binary outcomes associated with a large number of covariates and frequently adopted in practice as a quasi-maximum likelihood estimator (QMLE) for conventional binary choice models (BCMs). Logistic regression assumes the BCM with logistic errors, i.e., the logit model, although this is not explicitly recognized when used in machine learning contexts, which are generally astructural and not based on any fully specified model. The use of logistic regression or any other QMLE to estimate a BCM is generally invalid, yielding inconsistent estimates unless the error term in the BCM follows the same distribution as that employed in the QMLE. Furthermore, there are many other existing approaches to consistently estimate BCMs proposed by various authors. [^1] They include Powell, Stock, and Stoker (1989), Ichimura (1993), Klein and Spady (1993), Ahn, Ichimura, Powell, and Ruud (2018), Khan, Lan, Tamer, and Yao (2024), among others. Nevertheless, logistic regression is still used more widely in actual applications than any of these alternatives, perhaps due to its computational simplicity and the availability of software packages.
@@
In this paper, we study consistency of QMLE for BCMs. Ruud (1983) provided conditions under which QMLE for BCMs may asymptotically yield a slope coefficient that is a constant multiple of the true slope coefficient. However, he did not fully establish the slope consistency of QMLE, which requires the existence of a positive multiple of the slope coefficient identified as an interior maximizer of the population QMLE likelihood function over an appropriately restricted parameter space. He showed neither that a solution to the first-order conditions for the maximizer of the likelihood function exists within the restricted parameter space, nor that the multiplicative constant is strictly positive. There may be no interior solution over the restricted parameter set. The multiplicative constant may also be zero or even negative, which could lead to incorrect conclusions of no effect or a reverse effect, respectively. We fill this gap by providing a formal proof of slope consistency under the same set of conditions for any binary choice model identified as in Horowitz (1992) or Manski (1975, 1985).
@@
More specifically, we consider a binary outcome $Y$ given by the sign of $Y^{\ast}=V-U=\alpha_{0}+X^{\prime}\beta_{0}-U$, where $X$ is an $m$-dimensional covariate, and $V$ and $U$ are the scalar-valued systematic component and error term, respectively. For this BCM, we assume the index dependence given by $\mathcal{L}(U|X)=\mathcal{L}(U|V)$ with $\mathcal{L}$ denoting distribution and the linearity in expectation given by $\mathbb{E}(X|V)=aV+b$ with $a,b\in\mathbb{R}^{m}$. [^2] The index dependence is commonly imposed in the literature on consistent estimation of index models including all papers cited in Footnote 1. Ruud (1983) assumed the linearity in expectation, as well as the independence of $X$ and $U$, in which case the index dependence is also trivially satisfied. We establish slope consistency of QMLE with no other restrictive assumptions, except for necessary identification conditions and high-level technical conditions such as concavity and differentiability of the likelihood function which are commonly imposed to establish consistency of the maximum likelihood estimator (MLE). The linearity in expectation is certainly restrictive, but it holds, for instance, if $X$ has an elliptical distribution. Moreover, we may weight observations on $X$ appropriately so that the weighted observations can be regarded as a distribution satisfying the condition. See Ruud (1986) and Newey and Ruud (1994).
For binary outcomes generated by a BCM, QMLE is thus a viable option to use. In fact, logistic regression is used extensively in machine learning to analyze binary outcomes associated with a large number of covariates. This paper shows that it actually yields a consistent slope coefficient of the underlying binary choice model from which the binary outcomes are generated, provided that it satisfies the required index dependence and linearity in expectation conditions, in addition to other standard technical conditions. Needless to say, for a high-dimensional binary choice model, it is crucial to estimate the slope coefficient consistently. Once the slope coefficient is estimated consistently, it is relatively simple and straightforward to consistently estimate the intercept, which is one-dimensional.
@@
The rest of the paper is organized as follows. In Section 2, we introduce our model and background theory with assumptions that are needed to ensure identification of the model and regularity in maximum likelihood estimation. Section 3 provides our main results on slope consistency of the QMLE with required assumptions on index dependence and linearity in expectation. Section 4 concludes the paper. Mathematical proofs are in the Appendix.
@@ 2 Preliminaries
where $\operatorname*{sgn}$ is the sign function defined as $\operatorname*{sgn}\,(z)=\pm 1$ for $z\geq 0$ and $z<0$, respectively, $X$ is an $m$-dimensional vector of covariates, $\theta_{0}=(\alpha_{0},\beta_{0}^{\prime})^{\prime}$ is the true value of $(m+1)$-dimensional parameter vector $\theta=(\alpha,\beta^{\prime})^{\prime}$, and $U$ is the error term. [^3] We may define $Y=1\{Y^{\ast}\geq 0\}$ so that $Y$ takes value 1 or 0, following the usual convention for binary choice model. This will not change any of our subsequent discussions in the paper. In the paper, we define $Y$ to take value 1 or 1-1 to make our theory more directly applicable for the logistic regression used in machine learning. Throughout the paper, we let $\theta\in\Theta$, where $\Theta$ is a convex subset of $\mathbb{R}^{m+1}$ that is arbitrarily large. [^4] This is necessary because we identify $\theta_{0}$ only up to a scalar multiplication in the derivation of our main result. Our definition of $\Theta$ is essentially equivalent to set $\Theta=\mathbb{R}^{m+1}$. In the paper, however, we define $\Theta$ separately from $\mathbb{R}^{m+1}$ to make our subsequent assumptions more directly comparable to those existing in the literature.
Since we focus on the consistency of slope coefficient up to a positive scalar—referred to as slope consistency for simplicity—we impose conditions to ensure that $\theta_{0}$ is identified up to a positive scalar. Throughout the paper, we use $\mathcal{L}$ to denote the law or distribution.
@@
Assumptions 2.1 and 2.2 are imposed for the BCM in Horowitz (1992), modified from Manski (1975, 1985), which ensure that $\theta_{0}=(\alpha_{0},\beta_{0}^{\prime})^{\prime}$ is identified up to multiplication by a positive scalar with $\beta_{0}\neq 0$.
We consider the maximum likelihood estimation of a possibly misspecified model, referred to as quasi-maximum likelihood (QML) estimation, which assumes that $U$ is independent of $X$ and has distribution function $F$. The QML estimator $\hat{\theta}$ is defined as the maximum of
@@
over a parameter set $\Theta$. We define
@@
Assumption 2.3 is standard. Part (a) assumes the existence of a pseudo-true value given by QMLE, which is commonly imposed in the existing work of QMLE, see, e.g., White (1982), Gourieroux, Monfort, and Trognon (1984). Part (b) ensures that $Q(\theta)$ is well-defined and $\hat{Q}(\theta)\to_{p}Q(\theta)$ for any $\theta\in\Theta$. Finally, Part (c) guarantees that $\theta_{\ast}$ is unique.
@@
Lemma 2.1 follows immediately from Theorem 2.7 in Newey and McFadden (1994), since all conditions required in their theorem are trivially satisfied under our Assumption 2.3.
Lemma 2.1 shows that $\hat{\theta}$ has a probability limit $\theta_{\ast}$. Similar results are provided by White (1982) and Gourieroux, Monfort, and Trognon (1984). Unlike these works, we assume neither the compactness of $\Theta$ nor the existence of an envelope function for $|\log F(\alpha+x^{\prime}\beta)|$ and $|\log(1-F(\alpha+x^{\prime}\beta))|$ with finite moment. Instead, we assume convexity of $\Theta$ and impose concavity assumptions on $\log F$ and $\log(1-F)$.
@@
Given Assumption 2.3, Assumption 2.4 implies that $\theta_{\ast}$ is uniquely defined as the solution to the first order condition (FOC) given by the derivative of $Q$. Both Part (c) of Assumption 2.3 and Assumption 2.4 are satisfied for normal and logistic distribution functions, which are most commonly used in practice.
Subsequently, we let Assumption 2.4, as well as Assumption 2.3, hold. Under these assumptions, $Q$ is strictly concave and twice continuously differentiable. Therefore, if we define
@@
the probability limit $\theta_{\ast}=(\alpha_{\ast},\beta_{\ast}^{\prime})^{\prime}$ of $\hat{\theta}$ is defined as the solution to the first-order condition $\dot{Q}(\theta)=0$, where $\dot{Q}$ is the first-order derivative of $Q$ given by
@@ 3 Slope Consistency
$V=\alpha_{0}+X^{\prime}\beta_{0}$
in what follows. We introduce two important assumptions that are crucial for the slope consistency of QMLE.
@@
Assumption 3.1 requires that the error distribution depends on $X$ only through the index $V$. Such an index dependence condition for the error distribution is often used in the literature. It’s used in, e.g., Klein and Spady (1993) for the BCM, and Ichimura (1993) for a class of single index models, among many others. This condition is satisfied trivially if $U$ is independent of $X$, as is often assumed for the BCM.
@@
Assumption 3.2 requires that the conditional mean of $X$ given $V$ is a linear function of $V$, which will be referred to simply as the linearity in expectation condition. This condition is certainly restrictive, but it holds, for instance, if $X$ has an elliptical distribution. We may also weight observations on $X$ appropriately so that the distribution of the weighted observations can be regarded as satisfying this condition. See, e.g., Ruud (1986) and Newey and Ruud (1994). In place of original observations $\{x_{i}\}_{i=1}^{n}$ on $X$, they use the observations weighted by $w_{i}=\sigma(x_{i})/\hat{\tau}(x_{i})$ for $i=1,\ldots,n$, where $\sigma$ is the standard multivariate normal density satisfying Assumption 3.2, and $\hat{\tau}$ is a local kernel density estimator of $X$. Note that the weighted observations $\{w_{i}x_{i}\}_{i=1}^{n}$ may be regarded approximately as the observations drawn from $X$ having density $\sigma$. This was proposed by Ruud (1986) for QMLE, and subsequently used by Newey and Ruud (1994) for weighted least squares. [^5] Ruud (1986) studies a class of limited dependent variable models given by $Y=\pi(\alpha_{0}+X^{\prime}\beta_{0}-U)$ with known $\pi$, and Newey and Ruud (1994) consider the semiparametric single index model given by $\mathbb{E}(Y|X)=\pi(X^{\prime}\beta_{0})$ with unknown $\pi$.
For slope consistency of QMLE, we consider the QMLE over a restricted set of parameters $\theta=(\alpha,\beta^{\prime})^{\prime}$ specified as
with some $c,r\in\mathbb{R}$. The resulting estimator will be referred to simply as the restricted QMLE. Note that the restricted QMLE has only two parameters, while the unrestricted QMLE has $(m+1)$-parameters. Let $\dot{Q}(c,r)$ be defined as $\dot{Q}(\theta)$ in (5) with the $(m+1)$-dimensional parameter $\theta=(\alpha,\beta^{\prime})^{\prime}$ being restricted to the two dimensional parameter $(c,r)$ introduced in (6). Moreover, we let
@@
for $v\in\mathbb{R}$. Under Assumption 2.1, we have $\Pi(v)\geq 1/2$ if $v\geq 0$, and $\Pi(v)<1/2$ if $v<0$.
It follows that
$\dot{Q}(c,r)$ $=\mathbb{E}\left(1\{U\leq V\}\dot{\ell}_{+}(cV+r)-1\{U>V\}\dot{\ell}_{-}(cV+r)\right)\begin{pmatrix}1\\ X\end{pmatrix}$
$=\mathbb{E}\left(\Pi(V)\dot{\ell}_{+}(cV+r)-(1-\Pi(V))\dot{\ell}_{-}(cV+r)\right)\begin{pmatrix}1\\ \mathbb{E}(X|V)\end{pmatrix}$
$=\mathbb{E}\left(\Pi(V)\dot{\ell}_{+}(cV+r)-(1-\Pi(V))\dot{\ell}_{-}(cV+r)\right)\begin{pmatrix}1\\ aV+b\end{pmatrix}$, (7)
where we use the law of iterated expectation by taking conditional expectation on $V$ before we take expectation $\mathbb{E}$ to define $\dot{Q}(c,r)$, and apply the linearity in expectation for $\mathbb{E}(X|V)$ in Assumption 3.2. Consequently, we may deduce from (7) that the FOC for $(c,r)$ given by the system of $(m+1)$-equations $\dot{Q}(c,r)=0$ effectively reduces to the following system of two equations
$\dot{Q}_{\bullet}(c,r)$ $=\mathbb{E}\left(1\{U\leq V\}\dot{\ell}_{+}(cV+r)-1\{U>V\}\dot{\ell}_{-}(cV+r)\right)\begin{pmatrix}1\\ V\end{pmatrix}$
$=\mathbb{E}\left(\Pi(V)\dot{\ell}_{+}(cV+r)-(1-\Pi(V))\dot{\ell}_{-}(cV+r)\right)\begin{pmatrix}1\\ V\end{pmatrix}=0$ (8)
@@
As discussed, Assumption 2.3 ensures that the likelihood function for the unrestricted QMLE has a unique interior maximizer asymptotically. This, however, does not mean that the likelihood function for the restricted QMLE also has a unique interior maximizer asymptotically. Therefore, we cannot automatically assume that the solution to the FOC in (3) exists. In fact, the slope consistency of QMLE requires that the FOC has a solution $(c_{\ast},r_{\ast})$ with $c_{\ast}>0$. Note that any solution $(c_{\ast},r_{\ast})$ to the FOC is a maximizer of the unrestricted QMLE, and therefore, it is a unique solution under Assumptions 3.1 and 3.2. In fact, Ruud (1983) shows neither that the FOC has a solution $(c_{\ast},r_{\ast})$, nor that $c_{\ast}>0$. He simply assumes that the likelihood function for his restricted QMLE, which is essentially identical to ours, has the maximizer given by a solution to the FOC. Li and Duan (1989) consider a more general class of models and introduce an additional high-level assumption which is sufficient to ensure for binary choice model that the FOC has a solution $c_{\ast}\in\mathbb{R}$, but not necessarily with $c_{\ast}>0$. See also Powell (1994) for related discussions.
@@
Let Assumptions 2.1, 2.2, 2.3, 2.4, 3.1 and 3.2 hold. Then $\dot{Q}_{\bullet}(c,r)=0$ has a unique solution $(c_{\ast},r_{\ast})$ for $c_{\ast}>0$ and $r_{\ast}\in\mathbb{R}$. Moreover, if we let $\theta_{\ast}=(\alpha_{\ast},\beta_{\ast}^{\prime})^{\prime}$ with $\alpha_{\ast}=c_{\ast}\alpha_{0}+r_{\ast}$ and $\beta_{\ast}=c_{\ast}\beta_{0}$, then $\hat{\theta}\to_{p}\theta_{\ast}$ as $n\to\infty$.

### v3

@@ Abstract
Although QMLE is generally inconsistent, logistic regression relying on the binary choice model (BCM) with logistic errors is widely used, especially in machine learning contexts with many covariates. This paper revisits the slope consistency of QMLE for BCMs. Ruud (1983) introduced a set of conditions under which QMLE may yield a constant multiple of the slope coefficient of BCMs asymptotically. However, he did not fully establish the slope consistency of QMLE, which requires the existence of a positive multiple of the true slope that maximizes the population QMLE likelihood over an appropriately restricted parameter space. We close this gap by providing a formal proof of slope consistency under the same set of conditions for BCMs identified as in Manski (1975, 1985). Our result implies that, under suitable conditions, logistic regression yields a consistent estimate of the slope coefficient for BCMs.
@@ 1 Introduction
Logistic regression is widely used in empirical work and in machine learning to analyze binary outcomes, and is often applied as a quasi-maximum likelihood estimator (QMLE) for binary choice models (BCMs). When the error distribution in the underlying BCM is not logistic, the logit likelihood is misspecified, and the QMLE need not be consistent. Despite many alternative approaches for consistently estimating semiparametric BCMs, [^1] E.g., Powell et al. (1989), Ichimura (1993), Klein and Spady (1993), Ahn et al. (2018), Khan et al. (2024), among others. logistic regression remains common in applications, likely due to its computational simplicity and the availability of software packages.
@@
This paper studies slope consistency of QMLE for BCMs. Ruud (1983) provided conditions under which QMLE for BCMs may asymptotically yield a slope vector proportional to the true slope. However, he did not fully establish the slope consistency of QMLE, which requires the existence of a positive multiple of the true slope that maximizes the population QMLE likelihood over an appropriately restricted parameter space. Without a careful argument establishing the existence of such a positive multiple, the proportionality constant need not be well-defined and, even when defined, could in principle be zero or negative, leading to incorrect conclusions (no effect or reversed sign). We close this gap by providing a formal proof of slope consistency under essentially the same conditions as in Ruud (1983) for BCMs identified as in Manski (1975, 1985).
@@
In the paper, we consider a binary outcome $Y$ given by the sign of $Y^{\ast}=\alpha_{0}+X^{\prime}\beta_{0}-U$ with $X$ taking values in $\mathbb{R}^{m}$, and denote throughout by $\mathcal{L}$ the law or distribution. We impose (i) index dependence $\mathcal{L}(U|X)=\mathcal{L}(U|V)$ with $V=\alpha_{0}+X^{\prime}\beta_{0}$, and (ii) linearity in expectation $\mathbb{E}(X|V)=aV+b$ for $a,b\in\mathbb{R}^{m}$.  [^2] The index dependence condition is commonly imposed in the literature on consistent estimation of index models, including the papers cited in Footnote 1. Ruud (1983) assumed linearity in expectation and independence of $X$ and $U$, which implies index dependence. Under (i)–(ii), together with necessary identification and regularity conditions (e.g., concavity and differentiability of the population log-likelihood), we establish slope consistency of the QMLE. Linearity in expectation is restrictive but holds, for instance, when $X$ is elliptically distributed, and can also be achieved by appropriate weighting; see Ruud (1986) and Newey and Ruud (1994).
Our results suggest that logistic regression can be used as a slope-consistent QMLE for an underlying BCM, provided that the index dependence and linearity-in-expectation conditions hold. This may provide some theoretical justification for the widespread use of logistic regression in machine learning for analyzing binary outcomes, as well as the popularity of logit and probit models in applied work.
@@
The rest of the paper is organized as follows. Section 2 introduces the model and background theory with assumptions for identification and regularity in MLE. Section 3 provides the slope consistency of the QMLE. Section 4 concludes the paper. Mathematical proofs are in the Online Supplement.
@@ 2 Preliminaries
where $\operatorname*{sgn}$ is the sign function defined as $\operatorname*{sgn}\,(z)=\pm 1$ for $z\geq 0$ and $z<0$, respectively, $X$ is an $m$-dimensional vector of covariates, $\theta_{0}=(\alpha_{0},\beta_{0}^{\prime})^{\prime}$ is the true value of $\theta=(\alpha,\beta^{\prime})^{\prime}$, and $U$ is the error term. Let $\theta\in\Theta$, where $\Theta=\mathbb{R}^{m+1}$.
Since we focus on the consistency of slope coefficient up to a positive scalar—referred to as slope consistency for simplicity—we impose conditions to ensure that $\theta_{0}$ is identified up to a positive scalar.
@@
Assumptions 2.1 and 2.2 ensure that $\theta_{0}=(\alpha_{0},\beta_{0}^{\prime})^{\prime}$ is identified up to multiplication by a positive scalar with $\beta_{0}\neq 0$; See, e.g., Manski (1975, 1985).
We consider the maximum likelihood estimation of a possibly misspecified model, referred to as quasi-maximum likelihood (QML) estimation, which assumes that $U$ is independent of $X$ and has distribution function $F$. The QML estimator $\hat{\theta}=(\hat{\alpha},\hat{\beta}^{\prime})^{\prime}$ is defined as the maximum of
@@
over $\Theta$. We define
@@
Assumption 2.3 is standard. Part (a) assumes the existence of a pseudo-true value given by QMLE, which is commonly imposed in the existing work of QMLE, see, e.g., White (1982), Gourieroux et al. (1984). Part (b) ensures that $Q(\theta)$ is well-defined and $\hat{Q}(\theta)\to_{p}Q(\theta)$ for any $\theta\in\Theta$. Finally, Part (c) guarantees that $\theta_{\ast}$ is unique.
@@
Lemma 2.1 shows that $\hat{\theta}$ has a probability limit $\theta_{\ast}$. Lemma 2.1 follows immediately from Theorem 2.7 in Newey and McFadden (1994), since all conditions required in their theorem are trivially satisfied under our Assumption 2.3.
@@
Given Assumption 2.3, Assumption 2.4 implies that $\theta_{\ast}$ is uniquely defined as the solution to the first order condition (FOC) given by the derivative of $Q$. Both Assumption 2.3(c) and Assumption 2.4 are satisfied for normal and logistic distribution functions, which are most commonly used in practice.
Subsequently, we let Assumptions 2.3 and 2.4 hold. Then $Q$ is strictly concave and twice continuously differentiable. Therefore, if we define
@@
the probability limit $\theta_{\ast}=(\alpha_{\ast},\beta_{\ast}^{\prime})^{\prime}$ of $\hat{\theta}$ is defined as the solution to the FOC $\dot{Q}(\theta)=0$, where $\dot{Q}$ is the first-order derivative of $Q$ given by
@@ 3 Slope Consistency
$V=\alpha_{0}+X^{\prime}\beta_{0},$
and consider the QMLE over a restricted set of parameters $\theta=(\alpha,\beta^{\prime})^{\prime}$ specified as
with some $c,r\in\mathbb{R}$. The resulting estimator will be referred to as the restricted QMLE. Note that the restricted QMLE has only two parameters, while the unrestricted QMLE has $(m+1)$-parameters. To analyze the restricted QMLE, we rewrite $\dot{Q}(\theta)$ in (5) as
$\dot{Q}(c,r)=\mathbb{E}\left(1\{U\leq V\}\dot{\ell}_{+}(cV+r)-1\{U>V\}\dot{\ell}_{-}(cV+r)\right)\begin{pmatrix}1\\ X\end{pmatrix}$ (7)
with the $(m+1)$-dimensional parameter $\theta=(\alpha,\beta^{\prime})^{\prime}$ being restricted to the two dimensional parameter $(c,r)$ introduced in (6). Then we may easily deduce that
Lemma 3.1.
Let Assumptions 2.1, 2.2, 2.3 and 2.4 hold. Then the QMLE is slope consistent if and only if $\dot{Q}(c,r)=0$ has a solution $(c_{\ast},r_{\ast})$ such that $c_{\ast}>0$ and $r_{\ast}\in\mathbb{R}$.
Now we introduce a set of sufficient conditions to ensure that $\dot{Q}(c,r)=0$ has a solution $(c_{\ast},r_{\ast})$ such that $c_{\ast}>0$ and $r_{\ast}\in\mathbb{R}$.
@@
Assumption 3.1 requires that the error distribution depends on $X$ only through the index $V$. Such an index dependence condition for the error distribution is often used in the literature. [^3] See, e.g., Klein and Spady (1993) for the BCM, and Ichimura (1993) for a class of single index models.
@@
Assumption 3.2 requires that the conditional mean of $X$ given $V$ is a linear function of $V$, which will be referred to simply as the linearity in expectation condition. This condition is restrictive, but it holds, for instance, if $X$ has an elliptical distribution. We may also weight observations on $X$ appropriately so that the distribution of the weighted observations can be regarded as satisfying this condition; see, e.g., Ruud (1986) and Newey and Ruud (1994). [^4] Specifically, in place of the original observations $\{x_{i}\}_{i=1}^{n}$ on $X$, they use the observations weighted by $w_{i}=\sigma(x_{i})/\hat{\tau}(x_{i})$ for $i=1,\ldots,n$, where $\sigma$ is the standard multivariate normal density satisfying Assumption 3.2, and $\hat{\tau}$ is a local kernel density estimator of $X$. The weighted observations may be approximately regarded drawn from a distribution with density $\sigma$.
In what follows, under Assumptions 2.1, 2.2, 2.3 and 2.4, we establish that Assumptions 3.1 and 3.2 indeed ensure the existence of solution $(c_{\ast},r_{\ast})$ with $c_{\ast}>0$ to $\dot{Q}(c,r)=0$. It is clear that $\dot{Q}(c,r)=0$ may have such a solution without Assumptions 3.1 and 3.2, so they need not be necessary.
Let
@@
for $v\in\mathbb{R}$. Due to Assumption 2.1, we have $\Pi(v)\geq 1/2$ if $v\geq 0$, and $\Pi(v)<1/2$ if $v<0$. Under Assumptions 3.1 and 3.2, $\dot{Q}(c,r)$ defined in (7) becomes [^5] To see this, first use the law of iterated expectation (LIE) by taking the conditional expectation on $X$, where we simplify $\mathbb{E}(1\{U\leq V\}|X)=\Pi(V),\mathbb{E}(1\{U>V\}|X)=1-\Pi(V)$ under Assumption 3.1. Second, use LIE by taking the conditional expectation on $V$ and note $\mathbb{E}(X|V)=aV+b$ by Assumption 3.2.
$\dot{Q}(c,r)=\mathbb{E}\left(\Pi(V)\dot{\ell}_{+}(cV+r)-(1-\Pi(V))\dot{\ell}_{-}(cV+r)\right)\begin{pmatrix}1\\ aV+b\end{pmatrix}$, (8)
and consequently, the system of $(m+1)$-equations $\dot{Q}(c,r)=0$ effectively reduces to the system of two equations given by
$\dot{Q}_{\bullet}(c,r)$ $=\mathbb{E}\left(\Pi(V)\dot{\ell}_{+}(cV+r)-(1-\Pi(V))\dot{\ell}_{-}(cV+r)\right)\begin{pmatrix}1\\ V\end{pmatrix}=0$ (9)
@@
Slope consistency of the QMLE requires that the FOC $\dot{Q}_{\bullet}(c,r)=0$ has a solution $(c_{\ast},r_{\ast})$ with $c_{\ast}>0$. However, the existence of such a solution is not automatic. Even if Assumption 2.3 ensures that the population likelihood for the unrestricted QMLE has a unique maximizer, it does not mean that the population likelihood for the restricted QMLE also has a maximizer. Moreover, even if the FOC admits a solution, the associated $c_{\ast}$ need not be positive.
In fact, Ruud (1983) shows neither that the FOC exists a solution $(c_{\ast},r_{\ast})$ nor that $c_{\ast}>0$; instead, he assumes that the restricted likelihood (which is essentially identical to ours) is maximized at a point satisfying the FOC. Li and Duan (1989) consider a more general class of models and impose an additional high-level condition that guarantees, for BCMs, existence of a solution $c_{\ast}\in\mathbb{R}$, but not necessarily with $c_{\ast}>0$. See also Powell (1994) for related discussion. We address these issues by establishing the following lemma, which is the key technical contribution of the paper.
@@
The proofs for Lemma 3.2 and Theorem 3.3 are provided in Online Supplement.
@@
Let Assumptions 2.1, 2.2, 2.3, 2.4, 3.1 and 3.2 hold. Then $\dot{Q}_{\bullet}(c,r)=0$ has a unique solution $(c_{\ast},r_{\ast})$ for $c_{\ast}>0$ and $r_{\ast}\in\mathbb{R}$. Moreover, $\hat{\alpha}\to_{p}c_{\ast}\alpha_{0}+r_{\ast}$ and $\hat{\beta}\to_{p}c_{\ast}\beta_{0}$ as $n\to\infty$.
@@
Under suitable regularity conditions, $\sqrt{n}(\hat{\theta}-\theta_{\ast})=-\big[\ddot{Q}_{n}(\theta_{\ast})\big]^{-1}\big[\sqrt{n}\dot{Q}_{n}(\theta_{\ast})\big]+o_{p}(1)$ and $\sqrt{n}\dot{Q}_{n}(\theta_{\ast})$ has a normal limit distribution, where $\dot{Q}_{n}$ and $\ddot{Q}_{n}$ are the first and second derivatives, respectively, of $Q_{n}$ defined in (2). In this case, inference for $\beta_{\ast}$ in $\theta_{\ast}=(\alpha_{\ast},\beta_{\ast}^{\prime})^{\prime}$ can be conducted using standard QMLE theory with a robust (sandwich) variance; see White (1982). Since $\beta_{\ast}=c_{\ast}\beta_{0}$, we can test scale-invariant hypotheses about $\beta_{0}$, such as $\beta_{j,0}=0$ and $\beta_{j,0}=\beta_{k,0}$ for $j\neq k$, among many others. Such scale-invariant hypotheses are natural: in our setup $\beta_{0}$ is identified only up to a positive scale, and even in BCMs with a scale normalization, the normalization is typically a convention that fixes the scale of latent utility and carries no economic content.